# Spatial, Spectral and Temporal Response of High Intensity Laser Plasma Mirrors–Direct Observation of the Ponderomotive Push

### v2

@@ IV Simulation
To complement our experimental observations, we performed 3D3V and 2D3V Particle-In-Cell (PIC) simulations using the massively parallel, fully relativistic, open-source OSIRIS 4.0 framework [46]. These simulations provide a detailed understanding of plasma dynamics at resolutions and time scales inaccessible in experiments, offering insights into energy deposition, charge transport, and harmonic generation.
@@
The depth of the plasma surface depression due to the ponderomotive push exerted by the laser field was analyzed as a function of $a_{0}$, as shown in Figure 7 . The results indicate a monotonic increase in penetration depth with increasing laser intensity, consistent with theoretical predictions of ponderomotive scaling.
The modification of plasma density due to laser interaction is illustrated in Figure 8 , showing snapshots of electron, ion, and charge densities. The charge separation induced by the laser field results in the formation of a characteristic plasma cavity at higher intensities.
To analyze the impact of laser intensity and plasma density on the reflected electromagnetic field, we performed a Fourier analysis of the transverse field component. Notably, the reflected field exhibits odd harmonics, a signature of relativistic plasma boundary oscillations, which arise due to nonlinear plasma interactions[47, 48].
@@ V Conclusion
We have reported direct measurement of the spatial, spectral, and temporal characteristics of light reflected from relativistic PMs, in agreement with both particle-in-cell (PIC) simulations and theoretical models. We present a direct experimental measurement of the shape of the critical electron density surface in a high intensity laser excited plasma on a solid surface for two different laser wavelengths, providing clear evidence of the ponderomotive push. The behavior of PMs was investigated using high- and ultrahigh-contrast pulses. Significant spectral broadening and pulse shape modification are observed, arising from the deformation of the PM and various nonlinear plasma dynamics. As the conventional metal-coated or dielectric mirrors are unsuitable at relativistic intensities due to their low damage thresholds, relativistic PMs provide a promising alternative for reaching extreme intensities and exploring quantum electrodynamics (QED) regimes. The curvature induced to the reflected beam by the PM can be utilized in the far field for multiple applications, such as controlling the divergence of attosecond pulses generated during the interaction or focusing the reflected beam.

### v3

@@ IV Simulation
To complement our experimental observations, we performed 3D3V and 2D3V Particle-In-Cell (PIC) simulations using the massively parallel, fully relativistic, open-source OSIRIS 4.0 framework [fonseca2002osiris]. These simulations provide a detailed understanding of plasma dynamics at resolutions and time scales inaccessible in experiments, offering insights into energy deposition, charge transport, and harmonic generation.
@@
The depth of the plasma surface depression due to the ponderomotive push exerted by the laser field was analyzed as a function of $a_{0}$, as shown in Fig. 9 . The results indicate a monotonic increase in penetration depth with increasing laser intensity, consistent with theoretical predictions of ponderomotive scaling. The modification of plasma density due to laser interaction is illustrated in Fig. 10 , showing snapshots of electron, ion, and charge densities. The charge separation induced by the laser field results in the formation of a characteristic plasma cavity at higher intensities. To analyze the impact of laser intensity and plasma density on the reflected electromagnetic field, we performed a Fourier analysis of the transverse field component. Notably, the reflected field exhibits odd harmonics, a signature of relativistic plasma boundary oscillations, which arise due to nonlinear plasma interactions[Baeva2006, Ganeev2007].
@@ V Conclusion
We have reported direct measurement of the spatial, spectral, and temporal characteristics of light reflected from relativistic PMs, in agreement with both particle-in-cell (PIC) simulations and theoretical models. We present a direct experimental measurement of the shape of the critical electron density surface in different kind of high intensity laser excited plasma on a solid surface for two different laser wavelengths, providing clear evidence of the ponderomotive push. The average velocity of the deformation observed for Setup-A was $0.043c$ at the intensity of $3\times 10^{19}~W/\mathrm{cm}^{2}$, in line with the theoretical value of $0.025c$ at $1.5\times 10^{18}~W/\mathrm{cm}^{2}$ reported by Wilks et al. (1992) [wilks1992absorption].
The behavior of PMs was investigated using high- and extreme-contrast pulses. Significant spectral broadening and pulse shape modification are observed, arising from the deformation of the PM and various nonlinear plasma dynamics. As the conventional metal-coated or dielectric mirrors are unsuitable at relativistic intensities due to their low damage thresholds, relativistic PMs provide a promising alternative for reaching extreme intensities and exploring quantum electrodynamics (QED) regimes. The curvature induced to the reflected beam by the PM can be utilized in the far field for multiple applications, such as controlling the divergence of attosecond pulses generated during the interaction or focusing the reflected beam.
VI Acknowledgements
AD acknowledges support from the Anusandhan National Research Foundation (ANRF) of the Government of India through core grant CRG/2022/002782 as well as a J C Bose Fellowship grant JCB/2017/000055. GRK acknowledges major support for this research from the grant “Physics and Astronomy (Project Identification No. RTI4002) Department of Atomic Energy, Tata Institute of Fundamental Research” and partially from the grant JBR/2021/00039 of the Anusandhan National Research Foundation (ANRF), both of the Government of India.